OpenStudy (anonymous):

How many real solutions are there to the equation x2 − 7x + 12 = 0?

OpenStudy (anonymous):

There is two

OpenStudy (dangerousjesse):

Please don't give out direct answers :)

OpenStudy (shadowlegendx):

@Alysa007 Explain how you got that

OpenStudy (shadowlegendx):

x2 − 7x + 12 = 0? Use the formula: b^2 -4ac

OpenStudy (anonymous):

Sorry lol... And Shadow I did the math

OpenStudy (shadowlegendx):

1 = a -7 = b 12 = c

OpenStudy (shadowlegendx):

-7^2 = 49 49 - 4(1)(12) = 48 49 - 48 = 1

OpenStudy (shadowlegendx):

If you get 0, then there is only one solution. If you get any number besides 1, you have two solutions. Since 1 is the square root of 1, it is a real solution Two real solutions